# Dynamics of Two Species with Density-Dependent Interactions in a Mutualistic Context

### v1

@@ 4.3 From Parasitism to Mutualism: Parameter-Driven Transitions in Interaction Dynamics
In this section, we include models that, while not always strictly mutualistic, satisfy Definition 2.1, and fit to our general framework.
Depending on the position in the phase portrait and the population densities of each species, a model satisfying these assumptions could represent a parasitic interaction in one region, followed by a competitive interaction in another region and with different species densities, then a mutualistic one, and then return to a parasitic interaction at yet other densities. As Theorem 3.1 is stated in a more general context, it remains valid despite the specific transition dynamics examined in Section 2.2. Additionally, Corollaries 3.1 and 3.2 characterize the nature of equilibrium points with index +1+1 when intraspecific competition remains constant in sign. However, a given model representing a specific type of interaction can describe an entirely different interaction when its parameter values change (as for example intraspecific competition) while still remaining within the framework of Theorem 3.1.
In this section, we examine a first model which, depending on the signs of its coefficients—and thus its partial derivatives—can represent either a parasitic or a mutualistic interaction, as discussed in Section 2.2. We analyze the consequences of this change on phase portraits and equilibrium points of index +1+1. We then consider a second model that falls within Definition 2.1, where we keep the coefficient signs unchanged but slightly modify the values of certain parameters. We observe that an equilibrium previously identified as attractive in a parasitic region remains attractive but now in a mutualistic region for both species. The following examples are consistent with assumptions H1 to H4.
@@
$\begin{cases}\dot{x}=\delta x\left(a_{x}-b_{x}\left(x-c_{x}\right)^{2}-d_{x}y\right),\\ \dot{y}=\delta y\left(a_{y}-b_{y}\left(y-c_{y}\right)^{4}-d_{y}x\right),\end{cases}$ (8)
@@
Figure 8: Phase portraits of system (8) illustrating how changes in parameter signs alter the interaction type. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. With: $a_{x}=8$, $b_{x}=1$, $c_{x}=5.5$, $d_{x}=1$ and $a_{y}=9.5$, $b_{y}=0.5$, $c_{y}=4$, $d_{y}=1$.
@@
$\begin{cases}\dot{x}=x(-a_{x}-b_{x}(x-c_{x})^{2}+d_{x}y),\\ \dot{y}=y(a_{y}-b_{y}(x-c_{y})^{4}-d_{y}y).\end{cases}$ (9)
@@
(a) Phase portrait of (9), with parameters: $a_{x}=1.7$, $c_{x}=5.2$ and $c_{y}=5$.
(b) Phase portrait of (9), with parameters: $a_{x}=1.5$, $c_{x}=1.9$ and $c_{y}=7$.
Figure 9: Phase portraits illustrating how changing parameter values—while maintaining the interaction type—modifies the equilibrium locations. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. Black points represent attractive points and gray points represent saddle points. System (9) with: $b_{x}=0.4$, $d_{x}=1$, $a_{y}=6.5$, $b_{y}=0.03$ and $d_{y}=1$.
@@ 5 Discussion
In this study we investigated the deterministic dynamics of two interacting species under density-dependent interactions, with a particular focus on mutualistic interactions. Our approach allowed to draw a general theoretical framework, formalized by Theorem 3.1, which is independent of the specific type of ecological interaction considered, covering for instance predator-prey, competitive and mutualistic interactions. However, while most mutualistic models reviewed in Hale and Valdovinos, (2021) satisfy assumptions H1 to H4, some mutualistic models do not necessarily adhere to these assumptions. For instance, they may involve more than two isoclines intersecting at a single point, or include emigration effects (Thompson et al.,, 2006; Holland and DeAngelis,, 2010). Our framework nevertheless provides general predictions of the dynamics of species involved in a wide array of ecological interactions, and highlights the key-role of density-dependent processes, that are likely widespread in natural communities (Kawatsu and Kondoh,, 2018).
@@
We extended the classical definition of mutualism, allowing density-dependent transitions between mutualistic and antagonistic interactions. We thus proposed a broader Definition 2.1: a system is considered mutualistic if there exists at least one region for each species in the phase portrait where each benefits from the presence of the other. This relaxed definition enables the exploration of more diverse ecological dynamics and highlights biologically realistic features of species interactions. In textbook example of mutualistic interactions, such as pollination of flowering plants by insects, large densities of invasive bees have been documented to decrease plant reproductive success because of the damages caused by over-visitation of flowers (Aizen et al.,, 2014). As shown in some of the mutualistic models we studied, it is possible that small changes in parameter values could cause an equilibrium point that was initially located in a strictly mutualistic region to shift into a parasitic one. By allowing such transitions between mutualism and parasitism, our extended framework opened to the analysis of more complex systems and revealed new dynamical phenomena such as the emergence of limit cycles. Our theoretical predictions therefore provide a more realistic view of population dynamics of interacting species, opening new research avenues for community ecology.
@@ 6 Proofs
In this section, we provide the proofs of the different results stated in Section 3 and 4. We begin with the proof of Theorem 3.1, which describes the alternation of equilibrium point indices along the isoclines.
@@
Let $\left(x,y\right)$ an equilibrium point in the strictly positive quadrant. We can calculate the Jacobian of the system (1) at the point $\left(x,y\right)$:
@@
We end with the proof of Theorem 4.3 and 4.4 which provide conditions under which stable limit cycles may be observed.
Proof of Theorem 4.3.
We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 5. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
@@
Our goal now is to demonstrate that the limit set $\omega(\gamma)$ of the trajectory $\gamma$ is a limit cycle. According to the fivefold way (see Coleman, (1983), page 265), this will follow if we can show that $\omega(\gamma)$ contains no critical points. The critical points of system (1) are located at the origin $(0,0)$, at the point $(x_{2},0)$ on the $x$-axis, and at the point $\left(x^{*},y^{*}\right)$. $\left(x^{*},y^{*}\right)$ cannot be in $\omega(\gamma)$ because $\left(x^{*},y^{*}\right)$ is repusilve by assumption 1.9. Moreover we cannot have that $\omega(\gamma)=\{(0,0)\}$. In fact, for the trajectory $\gamma$ to approach $(0,0)$, it must intersect the interior of the region where $f>0$ and $g<0$ (as shown in Figure 5), and stay in this region as $t\to+\infty$. However, if $\gamma$ is in the interior of this region, then $f>0$ and $\dot{x}>0$, which implies that we cannot have both $x(t)$, $y(t)$ remaining in the region for all large enough $t$ and $(x(t),y(t))\rightarrow(0,0)$ as $t\rightarrow+\infty$. A similar argument shows that $\omega(\gamma)\neq\{(x_{2},0)\}$.
Thus the last case is if $\omega(\gamma)$ contains any critical points at all, $\omega(\gamma)$ can be a cycle graph. $\left(x^{*},y^{*}\right)$ cannot belong to a cycle graph since $\left(x^{*},y^{*}\right)$ is not the $\omega$-limit set of any nonconstant trajectory. If $(x_{2},0)$ belongs to a cycle graph, then so must the orbit along $B$ with $\omega$-limit set $\{(x_{2},0)\}$ and $\alpha$-limit set $\{(0,0)\}$, since that is the only orbit in the rectangular region with $\omega$-limit set $\{(x_{2},0)\}$. But in a cycle graph, each limit set of each orbit is a critical point of the cycle graph. We must then have that $\{(0,0)\}$ in the cycle graph, then the unbounded orbit on $A$ also belongs to the cycle graph since it is the only orbit in the positive quadrant whose $\omega$-limit set is $\{(0,0)\}$. This is impossible since $\omega(\gamma)$ is bounded because $\gamma$ is bounded. Thus $\omega(\gamma)$ contains no critical points. By the fivefold way (see Coleman, (1983), page 265) we have that $\omega(\gamma)$ is a cycle. ∎
Proof of Theorem 4.4.
We will follow the same approach as in the previous proof. We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 7. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.

### v2

@@ 4.3 From Parasitism to Mutualism: Parameter-Driven Transitions in Interaction Dynamics
In this section, we illustrate how our framework can be applied to models that cannot be captured by a definition of strict mutualism. We propose two models that, while not always strictly mutualistic, satisfy Definition 2.1, and fit to our general framework. Depending on population densities and on the location of the equilibrium in the phase portrait, such models may represent parasitic, competitive, or mutualistic interactions in different regions. In particular, Definition 2.1 allows for systems in which the nature of species interactions may change continuously as parameters vary, while still remaining within the same mathematical structure. Importantly, the general results established in Theorem 3.1 apply independently of transitions between mutualism and parasitism, as the theorem is formulated under assumptions that are broader than those imposed in Section 2.2. As a consequence, a model within this framework may describe qualitatively different ecological interactions solely through continuous changes in parameter values—such as the strength or sign of intraspecific competition—without leaving the scope of Theorem 3.1.
First, we examine a model which, depending on the signs of its coefficients—and thus its partial derivatives—can represent either a parasitic or a mutualistic interaction, as discussed in Section 2.2. We analyze the consequences of this change on phase portraits and equilibrium points of index +1+1. We then consider a second model that falls within Definition 2.1, where we keep the coefficient signs unchanged but slightly modify the values of certain parameters. We observe that an equilibrium previously identified as attractive in a parasitic region remains attractive but now in a mutualistic region for both species. The following examples are consistent with assumptions H1 to H4.
@@
$\begin{cases}\dot{x}=\delta x\left(a_{1}-b_{1}\left(x-c_{1}\right)^{2}-d_{1}y\right),\\ \dot{y}=\delta y\left(a_{2}-b_{2}\left(y-c_{2}\right)^{4}-d_{2}x\right),\end{cases}$ (4)
@@
Figure 8: Phase portraits of system (4) illustrating how changes in parameter signs alter the interaction type. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. With: $a_{1}=8$, $b_{1}=1$, $c_{1}=5.5$, $d_{1}=1$ and $a_{2}=9.5$, $b_{2}=0.5$, $c_{2}=4$, $d_{2}=1$.
@@
$\begin{cases}\dot{x}=x(-a_{1}-b_{1}(x-c_{1})^{2}+d_{1}y),\\ \dot{y}=y(a_{2}-b_{2}(x-c_{2})^{4}-d_{2}y).\end{cases}$ (5)
@@
(a) Phase portrait of (5), with parameters: $a_{1}=1.7$, $c_{1}=5.2$ and $c_{2}=5$.
(b) Phase portrait of (5), with parameters: $a_{1}=1.5$, $c_{1}=1.9$ and $c_{2}=7$.
Figure 9: Phase portraits illustrating how changing parameter values—while maintaining the interaction type—modifies the equilibrium locations. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. Black points represent attractive points and gray points represent saddle points. System (5) with: $b_{1}=0.4$, $d_{1}=1$, $a_{2}=6.5$, $b_{2}=0.03$ and $d_{2}=1$.
@@ 5 Discussion
In this study we investigated the deterministic dynamics of two interacting species under density-dependent interactions, with a particular focus on mutualistic interactions. Our approach allowed to draw a general theoretical framework, formalized by Theorem 3.1, which is independent of the specific type of ecological interaction considered, covering for instance predator-prey, competitive and a wide range of mutualistic interactions, as highlighted in Tables 1, 2 and 3. However, while most mutualistic models reviewed in Hale and Valdovinos (2021) satisfy assumptions H1 to H4, some mutualistic models do not necessarily adhere to these assumptions: for instance, they may involve more than two isoclines intersecting at a single point, or include emigration (Thompson et al., 2006; Holland and DeAngelis, 2010).
@@
Importantly, we extended the classical definition of mutualism, allowing density-dependent transitions between mutualistic and antagonistic interactions. Such transition in the type of ecological interactions are well documented at the evolutionary scale (see for instance Drew et al. (2021)), and could be favored in ecological context where relative densities among interacting species might change. We thus proposed a broader Definition 2.1: a system is considered mutualistic if there exists at least one region for each species in the phase portrait where each benefits from the presence of the other. This relaxed definition enables the exploration of more diverse ecological dynamics and highlights biologically realistic features of species interactions, that can crucially depend on the relative densities of interacting species (see for instance density-dependent mutualistic effects in ant-aphid interactions, Breton and Addicott (1992)). Even in the textbook examples of mutualistic interactions, such as pollination of flowering plants by insects, large densities of invasive bees have been documented to decrease plant reproductive success because of the damages caused by over-visitation of flowers (Aizen et al., 2014). As shown in some of the mutualistic models we studied, it is possible that small changes in parameter values could cause an equilibrium point that was initially located in a strictly mutualistic region to shift into a parasitic one. By allowing such transitions between mutualism and parasitism, our extended framework sheds new light on the qualitative structure of species interactions and permits the emergence of dynamical phenomena not accessible under strictly mutualistic assumptions, including the theoretical possibility of limit cycles.
Numerous empirical studies report cyclic population dynamics in nature, which are overwhelmingly interpreted as the outcome of predator–prey or parasitic interactions (Turchin, 2003, p. 383-384). We showed that cycles can indeed not occur in case of strict mutualism. However, we also showed that periodic oscillations can occur when there are transitions between mutualistic and parasitic interaction regimes. To our knowledge, three contexts can give rise to periodic oscillations in a mutualism model: the two general contexts we give in Section 4.2 as well as the model analyzed by Song et al. (2025). Note that Song et al. (2025)’s model is a particular case of System 1 but it does not enter into the general context of Section 4.2, thus suggesting that periodic oscillations can occur in other situations. It is striking however that our examples and Song et al. (2025)’s model share a common feature: an Allee effect, i.e. a positive intraspecific feedback interactions when density is low. This observation raises the question of whether, beyond mutualism–parasitism transitions, an Allee effect may also be a necessary condition for the emergence of such cycles. Overall, our results propose an alternative and complementary interpretative framework for explaining some of the cyclic patterns observed in natural systems.
To our knowledge there is no empirical evidence of mutualistic interactions exhibiting sustained population cycles, but long-term survey of mutualistic populations with precise estimations of densities are actually scarce. Given the current lack of precise empirical data on the temporal dynamics of mutualistic populations, our results should primarily be interpreted as theoretical possibilities, which might nonetheless become relevant under particular ecological or evolutionary scenarios yet to be documented. Our framework thus highlights potential behaviors that could guide future empirical investigations aimed at identifying conditions under which mutualistic interactions may fluctuate rather than converge to equilibrium.
Finally, our framework provides general predictions of the dynamics of species involved in a wide array of ecological interactions, and highlights the key-role of density-dependent processes, that are widespread in natural communities (Kawatsu and Kondoh, 2018). In particular, the framework shows how density dependence can shape interaction outcomes across a continuum of ecological contexts. It does not rely on specific algebraic expressions for interaction terms but only on qualitative assumptions on the signs of partial derivatives, allowing it to encompass a broad family of models without requiring explicit functional forms. This flexibility opens the way to future extensions incorporating environmental variability, state-dependent interactions, and slow evolutionary changes in parameters.
@@ 6 Proofs
In this section, we provide the proofs of the different results stated in Sections 3 and 4. We begin with the proof of Theorem 3.1 and Proposition 3.5, which describe the alternation of equilibrium point indices along the isoclines.
@@
Proof of Proposition 3.5.
We restrict the analysis to the case where the equilibrium point lies on the $x$-axis i.e. is generated by an intersection of the isocline $\Gamma_{f}$ with the axis $y=0$. The case of an equilibrium point located on the $y$-axis, arising from an intersection of $\Gamma_{g}$ with the axis $x=0$, follows by symmetry.
Let $(x_{1},0)$ denote the equilibrium point closest to the origin $(0,0)$. Under our assumptions, there may exist at most one additional equilibrium point on the $x$-axis, denoted by $(x_{2},0)$, with $x_{1}<x_{2}$, which may or may not exist. We keep the same notations and conventions as in the proof of Theorem 3.1.
In a neighborhood of the equilibrium point $(x_{1},0)$, four local configurations of the vector field may occur, depending on the associated sign of the vector field across the adjacent regions. The equilibrium point $(x_{1},0)$ may have index +1+1, corresponding to an attractive or repulsive equilibrium. In this case, the alternation of signs across the four regions is either (+,),(,)(+,-),(-,-) or (,+),(+,+)(-,+),(+,+) illustrated in Figure 10(10(a)). Alternatively, $(x_{1},0)$ may have index 1-1, in which case it is a saddle point, and the alternation of signs is either (+,+),(,+)(+,+),(-,+) or (,),(+,)(-,-),(+,-) illustrated in Figure 10(10(b)). As in the interior case, the ordering of signs is read clockwise.
(a) Local possible structure of the phase portrait around an index +1+1 equilibrium on the $x$-axis.
(b) Local possible structure of the phase portrait around an index 1-1 equilibrium on the $x$-axis.
Figure 10: Four possible local configurations of the phase portrait near an equilibrium point located on the $x$-axis., Two possible configurations are shown for an equilibrium point with index +1+1 (10(a)) and index 1-1 (10(b)). The alternation of signs of the vector field across adjacent regions is read clockwise. These configurations correspond to equilibrium points generated by intersections of the isocline $\Gamma_{f}$ with the axis $x=0$.
Thus, when following the isocline $\Gamma_{f}$ and encountering an equilibrium point located in the strictly positive quadrant, the first intersection of $\Gamma_{f}$ with $\Gamma_{g}$ reached when moving along $\Gamma_{f}$ starting from $(x_{1},0)$, the alternation of signs is reversed, while still being read clockwise. More precisely, if the equilibrium point $(x_{1},0)$ on the $x$-axis is characterized by the alternation (+,),(,)(+,-),(-,-) in Figure 10(a), then the first equilibrium point encountered in the strictly positive quadrant exhibits the alternation $(+,+)$, (,+)(-,+), (,)(-,-), and (+,)(+,-), corresponding to case $(ii)$ in the proof of Theorem 3.1, which yields an equilibrium point with index 1-1. The remaining cases follow the same principle. The remaining alternations in the strictly positive quadrant then follow from Theorem 3.1.
It remains to determine the nature of the equilibrium point $(x_{2},0)$, when it exists. At this stage, we have shown that the index alternation persists when the equilibrium point $(x_{1},0)$ on the $x$-axis is taken into account. It remains to determine whether this alternation is also satisfied by a second equilibrium point on the axis, if such a point exists. This depends on whether the isocline $\Gamma_{g}$ intersects the $x$-axis, and on the number of such intersections. Indeed, even if intersections of $\Gamma_{g}$ with the $x$-axis do not generate additional equilibrium points, they modify the sign of the vector field in the vertical direction (for example changing a $(+,+)$ into a (+,)(+,-)).
If $\Gamma_{g}$ does not intersect the $x$-axis between $(x_{1},0)$ and $(x_{2},0)$, or if $\Gamma_{g}$ intersects it twice on this interval (resulting in a double sign change), then the alternation at $(x_{2},0)$ is the reverse of that at $(x_{1},0)$. For instance, if the alternation at $(x_{1},0)$ is $(+,+)$ and (,+)(-,+) in Figure 10(b), then at $(x_{2},0)$ it becomes (,+)(-,+) and $(+,+)$, which transforms an equilibrium point with index 1-1 into one with index +1+1. It remains to verify whether this behavior is consistent with the previously established index alternation. In both situations (zero or two intersections of $\Gamma_{g}$ with $x$-axis), $\Gamma_{g}$ may generate equilibrium points in the phase portrait through its intersections with $\Gamma_{f}$. Since the isoclines are assumed to be continuous and by Theorem 3.1, the number of equilibrium points created by intersections of $\Gamma_{f}$ and $\Gamma_{g}$ before $\Gamma_{f}$ reaches $(x_{2},0)$ is necessarily even. Consequently, the last equilibrium point encountered in the strictly positive quadrant, corresponding to the final intersection of $\Gamma_{f}$ and $\Gamma_{g}$, has the same index as $(x_{1},0)$. As shown above, the equilibrium point $(x_{2},0)$ has an index opposite to that of $(x_{1},0)$. Therefore, the alternation of equilibrium point indices is preserved when $\Gamma_{g}$ has zero or two intersections with the $x$-axis between $(x_{1},0)$ and $(x_{2},0)$.
It remains to consider the case where the isocline $\Gamma_{g}$ intersects the $x$-axis exactly once between $(x_{1},0)$ and $(x_{2},0)$. By the same reasoning as above, for an equilibrium point $(x_{1},0)$ characterized by a given alternation of sign pairs, the intersection of $\Gamma_{g}$ with the $x$-axis induces a change in the sign of the vector field in the $y$-direction along the axis. Then, when reaching $(x_{2},0)$, the intersection with $\Gamma_{f}$ induces a change in the sign of the vector field in the $x$-direction. For instance, if the alternation around $(x_{1},0)$ is given by $(+,+)$ and (,+)(-,+), then after crossing $\Gamma_{g}$ the alternation becomes (,)(-,-), and after crossing $\Gamma_{f}$ it becomes (+,)(+,-) at $(x_{2},0)$. As a result, the equilibrium point $(x_{2},0)$ has the same index as $(x_{1},0)$. Moreover, since $\Gamma_{g}$ is continuous, it intersects the isocline $\Gamma_{f}$ at least once between $(x_{1},0)$ and $(x_{2},0)$, and in fact an odd number of times. Consequently, the last equilibrium point generated by an intersection of $\Gamma_{f}$ and $\Gamma_{g}$ has an index opposite to that of $(x_{1},0)$, and hence opposite to that of $(x_{2},0)$. Therefore, the alternation of equilibrium point indices is preserved when following $\Gamma_{f}$ from $(x_{1},0)$ to $(x_{2},0)$.
@@
Let $\left(x,y\right)$ be an equilibrium point in the strictly positive quadrant. We can calculate the Jacobian of the system (1) at the point $\left(x,y\right)$:
@@
We now prove Corollary 3.8. Notice that the proof of Corollary 3.10 will be omitted since it follows exactly the same arguments.
Proof.
We suppose that $x_{1}>0$ and $(x_{1},0)$ is an equilibrium point on the $x$-axis, that is, $f(x_{1},0)=0$. We prove the result for equilibria on the $x$-axis, i.e. for equilibrium points given by the intersection of $\Gamma_{f}$ with the line $y=0$. The proof for equilibria on the $y$-axis, corresponding to the intersection of $\Gamma_{g}$ with the line $x=0$, is identical. We recall that $f\in\operatorname{\mathcal{C}^{1}}$ and that we assume $\frac{\partial f}{\partial x}(x,y)<0\quad\text{for all }(x,y)$.
We assume by contradiction that there exist two distinct equilibria $0<x_{1}<x_{2}$ on the $x$-axis such that $f(x_{1},0)=f(x_{2},0)=0$. Since $f$ is $C^{1}$, for $h>0$ a first-order Taylor expansion of $f(\cdot,0)$ at $x_{1}$ yields
$f(x_{1}+h,0)=f(x_{1},0)+h\frac{\partial f}{\partial x}(x_{1},0)+o\mathopen{}\left(h\right)=h\frac{\partial f}{\partial x}(x_{1},0)+o\mathopen{}\left(h\right),$
and
$f(x_{1}-h,0)=-h\frac{\partial f}{\partial x}(x_{1},0)+o\mathopen{}\left(h\right).$
Since $\frac{\partial f}{\partial x}(x_{1},0)<0$, we have $f(x,0)<0$ for $x>x_{1}$ sufficiently close to $x_{1}$, and $f(x,0)>0$ for $x<x_{1}$ sufficiently close to $x_{1}$. Similarly, a Taylor expansion at $x_{2}$ implies that $f(x,0)>0$ for $x<x_{2}$ sufficiently close to $x_{2}$. By continuity of $f(\cdot,0)$, there must then exist $x\in(x_{1},x_{2})$ such that $f(x,0)=0$, which contradicts hypotheses H1 to H4, which allow at most two equilibrium points on the axes. Hence, the equilibrium on the $x$-axis is unique.
Let $x_{1}$ be the unique equilibrium on the $x$-axis. From the previous sign analysis, $f(x,0)>0$ for $x<x_{1}$ close to $x_{1}$, and $f(x,0)<0$ for $x>x_{1}$ close to $x_{1}$. By hypothesis, the index of $(x_{1},0)$ is +1+1, and therefore, by Proposition 3.5, $(x_{1},0)$ is an attractive equilibrium point. ∎
@@
We present the proof of Proposition 4.6, which provides conditions under which stable limit cycles may arise.
Proof of Proposition 4.6.
The proof is based on the Poincaré-Bendixson theorem (Ciesielski, 2012), used here through the fivefold way formulation introduced by Coleman (1983). Our strategy consists in constructing a positively invariant region of the phase plane and analyzing the possible $\omega$-limit sets of a trajectory entering this region.
For the first case, we construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 5. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
@@
Our goal now is to demonstrate that the limit set $\omega(\gamma)$ of the trajectory $\gamma$ is a limit cycle. According to the Poincaré-Bendixson theorem, stated in the form of the fivefold way (see Coleman (1983), page 265), this will follow if we can show that $\omega(\gamma)$ contains no critical points. The critical points of system (1) are located at the origin $(0,0)$, at the point $(x_{2},0)$ on the $x$-axis, and at the point $\left(x^{*},y^{*}\right)$. $\left(x^{*},y^{*}\right)$ cannot be in $\omega(\gamma)$ because $\left(x^{*},y^{*}\right)$ is repulsive by assumption 1.9. Moreover we cannot have that $\omega(\gamma)=\{(0,0)\}$. In fact, for the trajectory $\gamma$ to approach $(0,0)$, it must intersect the interior of the region where $f>0$ and $g<0$ (as shown in Figure 5), and stay in this region as $t\to+\infty$. However, if $\gamma$ is in the interior of this region, then $f>0$ and $\dot{x}>0$, which implies that we cannot have both $x(t)$, $y(t)$ remaining in the region for all large enough $t$ and $(x(t),y(t))\rightarrow(0,0)$ as $t\rightarrow+\infty$. A similar argument shows that $\omega(\gamma)\neq\{(x_{2},0)\}$.
Thus the last case is if $\omega(\gamma)$ contains any critical points at all, $\omega(\gamma)$ can be a cycle graph. $\left(x^{*},y^{*}\right)$ cannot belong to a cycle graph since $\left(x^{*},y^{*}\right)$ is not the $\omega$-limit set of any nonconstant trajectory. If $(x_{2},0)$ belongs to a cycle graph, then so must the orbit along $B$ with $\omega$-limit set $\{(x_{2},0)\}$ and $\alpha$-limit set $\{(0,0)\}$, since that is the only orbit in the rectangular region with $\omega$-limit set $\{(x_{2},0)\}$. But in a cycle graph, each limit set of each orbit is a critical point of the cycle graph. We must then have that $\{(0,0)\}$ in the cycle graph, then the unbounded orbit on $A$ also belongs to the cycle graph since it is the only orbit in the positive quadrant whose $\omega$-limit set is $\{(0,0)\}$. This is impossible since $\omega(\gamma)$ is bounded because $\gamma$ is bounded. Thus $\omega(\gamma)$ contains no critical points. By the fivefold way (see Coleman (1983), page 265) we have that $\omega(\gamma)$ is a cycle. By the Poincaré-Bendixson theorem, applied through the fivefold way (see Coleman (1983), page 265) we have that $\omega(\gamma)$ is a cycle.
We will follow the same approach for the second case. We construct a rectangular region composed of sides $A$, $B$, $C$, and $D$, as shown in Figure 6. We will demonstrate that once a trajectory $\gamma$ enters this region, it cannot exit.
@@
We conclude with the proof of Proposition 4.8, which provides a necessary condition for the existence of a limit cycle.
Proof.
We consider system (1) on the strictly positive quadrant:
$\left\{\begin{array}[]{ll}\dot{x}=xf(x,y),\\ \dot{y}=yg(x,y).\end{array}\right$.
We assume that $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$ do not change sign.
We consider the Dulac function
$B(x,y)=\frac{1}{xy},$
which is $C^{1}$ on the strictly positive quadrant.
Let F(x,y)=(xf(x,y),yg(x,y))F(x,y)=(xf(x,y),\,yg(x,y)) denote the vector field. Then
$BF=\left(\frac{f(x,y)}{y},\,\frac{g(x,y)}{x}\right).$
Its divergence is
$\nabla\cdot(BF)=\frac{\partial}{\partial x}\left(\frac{f(x,y)}{y}\right)+\frac{\partial}{\partial y}\left(\frac{g(x,y)}{x}\right)=\frac{1}{y}\frac{\partial f}{\partial x}+\frac{1}{x}\frac{\partial g}{\partial y}.$
Since $x>0$ and $y>0$, the sign of $\nabla\cdot(BF)$ is entirely determined by the signs of $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$. If both $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$ are negative everywhere (or both postive everywhere), then $\nabla\cdot(BF)$ does not change sign and is not identically zero.
By the Bendixson–Dulac criterion (see Perko (2013)), a $C^{1}$ planar vector field defined on on the strictly positive quadrant cannot admit a periodic orbit entirely contained in that domain if there exists a $C^{1}$ Dulac function whose weighted divergence does not change sign.
Therefore, the system admits no limit cycle in the strictly positive quadrant. ∎